OpenStudy (anonymous):

Johnny has been planting blueberries in a garden in his backyard. Johnny started with blueberries in one square foot of garden modeled by the function s(x) = 10. He is allowed to convert more of the garden, and each additional square foot he maintains allows his blueberry plants to produce more by a rate of a(x) = 1.4x-2. Explain to Johnny how to create an equation to predict the number of blueberries he can expect based on the number of square feet he maintains. Describe how to determine the number of blueberries he will grow with 5 square feet. Use complete sentences. I have no idea

ganeshie8 (ganeshie8):

@ddriggers8

ganeshie8 (ganeshie8):

You can combine both the functions by multiplying them to create a new function that predicts the number of blueberries he can expect based on the number of square feet he maintains. \(\large t(x) = s(x) * a(x)\) \(\large t(x) = 10*1.4^{x-2}\)

ganeshie8 (ganeshie8):

plugin \(x=5\) in above equation to find the number of blueberries he will grow in 5 square feet land

ganeshie8 (ganeshie8):

\(\large t(5) = 10*1.4^{5-2}\) use ur calculator

OpenStudy (anonymous):

HOLY COW youre a life saver lol thank you SO much. i have just one more question I need help with, could you help me

ganeshie8 (ganeshie8):

sure, il try post it in a separate thread :)